DVICONCAT(1) DVICONCAT(1) NAME dviconcat−concatenateDVIfiles SYNOPSIS dviconcat[−mmagnification][−ooutfile]listoffiles DESCRIPTION Dviconcatconcatenatesthenamedfiles(orstandardinput,ifnofilenamesaregiven)ontothe standardoutputorthegivenoutfile,creatinganewDVIfileusablebyanyoftheTeXconversion programs,orevenbydviconcatitself. Themagnificationoftheoutputfileisthesameasthatofthefirstinputfileunlessthe−m optionisgiven.Whendviconcatchangesthemagnificationofanyinputfile,itprintsawarningto thiseffect. AUTHOR ChrisTorek,UniversityofMaryland SEEALSO dviselect(1),latex(1),tex(1) MC-TeXUser’sGuide BUGS Theprogramshouldbecalleddvicat,butthatnamewastaken(forthedriverfortheC/A/T typesetter).